OpenStudy (anonymous):

Need guidance please. I have an integral (sqrt(x^2+81)/2 dx. I know the 2 on the bottom is a constant but not sure how to start this prob,

OpenStudy (campbell_st):

\[1/2\int\limits (x^2 +81)^{1/2} dx\]

OpenStudy (anonymous):

then do I use the power rule for this?

OpenStudy (anonymous):

Is this what it should look like? \[1/2\int\limits (x^2 +81 )^ (3/2) dx x( 2x)\]

OpenStudy (anonymous):

oops meant 3/2 in front not 1/2

OpenStudy (campbell_st):

I think you will need a substitution

OpenStudy (anonymous):

yes that is what I am suppose to do but didn't only see u being (x^2+81) and du being 2x not sure what v would be for this question.

OpenStudy (campbell_st):

have you done hyperbolic functions...? as its \[1/2 \int\limits \sqrt{x^2 + 81} dx = \sinh(x/9)\]

OpenStudy (anonymous):

nope not yet. 3rd term calculus is what I am in. Integration by parts is what we learned most recent.

OpenStudy (anonymous):

Does anyone know how to do this by integration by parts?